# HiQ-Lip: The First Quantum-Classical Hierarchical Method for Global Lipschitz Constant Estimation of ReLU Networks

### v1

@@ Abstract
Estimating the global Lipschitz constant of neural networks is crucial for understanding and improving their robustness and generalization capabilities. However, precise calculations are NP-hard, and current semidefinite programming (SDP) methods face challenges such as high memory usage and slow processing speeds. In this paper, we propose HiQ-Lip, a hybrid quantum-classical hierarchical method that leverages Coherent Ising Machines (CIMs) to estimate the global Lipschitz constant. We tackle the estimation by converting it into a Quadratic Unconstrained Binary Optimization (QUBO) problem and implement a multilevel graph coarsening and refinement strategy to adapt to the constraints of contemporary quantum hardware. Our experimental evaluations on fully connected neural networks demonstrate that HiQ-Lip not only provides estimates comparable to state-of-the-art methods but also significantly accelerates the computation process. In specific tests involving two-layer neural networks with 256 hidden neurons, HiQ-Lip doubles the solving speed and offers more accurate upper bounds than the existing best method, LiPopt. These findings highlight the promising utility of small-scale quantum devices in advancing the estimation of neural network robustness.
@@ 1 Introduction
Neural networks have achieved remarkable success in various fields such as computer vision, natural language processing, and autonomous driving, establishing themselves as core technologies in modern artificial intelligence  (Zhao et al., 2024; Forner & Ozcan, 2023; Paniego et al., 2023). Despite these advancements, the robustness and generalization capabilities of neural networks remain active areas of research  (Djolonga et al., 2021; Bennouna et al., 2023). The global Lipschitz constant is a critical metric for measuring the robustness of a neural network’s output to input perturbations, playing a significant role in understanding and enhancing model robustness  (Leino et al., 2021a).
@@
However, accurately computing the global Lipschitz constant is an NP-hard problem, particularly for deep networks with nonlinear activation functions like ReLU  (Jordan & Dimakis, 2020). Due to this computational challenge, researchers have developed various approximation methods to estimate upper bounds of the global Lipschitz constant. Among these, the Formal Global Lipschitz constant (FGL) assumes that all activation patterns in the hidden layers are independent and possible, thereby providing an upper bound for the exact global Lipschitz constant  (Szegedy et al., 2014; Virmaux & Scaman, 2018).
Existing methods often rely on relaxations and semidefinite programming (SDP) to estimate the upper bound of the Lipschitz constant  (Chen et al., 2020; Shi et al., 2022). Although these methods have a solid theoretical foundation, they suffer from high memory consumption and slow computation speeds. Additionally, simpler approaches such as matrix norm approximations, while computationally efficient, tend to provide overly conservative bounds that fail to reflect the network’s true robustness accurately. These limitations drive the need for more efficient and precise methods to estimate the global Lipschitz constant of neural networks.
@@
In recent years, advancements in quantum computing have offered new avenues for tackling NP-hard problems  (Khumalo et al., 2022; Choi, 2008; Chatterjee et al., 2024). Given the computational complexity of global Lipschitz constant estimation, quantum computation is seen as a potential solution. Specifically, quantum devices based on the Quadratic Unconstrained Binary Optimization (QUBO) model, such as Coherent Ising Machines (CIMs) or quantum annealer, have demonstrated unique potential in solving complex combinatorial optimization problems  (Inagaki et al., 2016; Date et al., 2019). However, the limited number of qubits in current quantum computers poses significant challenges when directly applying them to neural network robustness evaluations. Although numerous works have attempted to address large-scale QUBO problems by employing strategies like divide-and-conquer and QUBO formula simplification to decompose them into smaller QUBO subproblems for quantum solving  (Pelofske et al., 2021; Zhou et al., 2023), applications in neural network robustness assessment remain largely unexplored.
@@
To bridge this gap and harness the acceleration capabilities of quantum computing for FGL estimation of neural network, we propose a quantum-classical hybrid hierarchical solving method named HiQ-Lip. HiQ means multilevel solution method for quantum classical mixing, while Lip stands for global Lipschitz constant. By reformulating the global Lipschitz constant estimation problem as a cut-norm problem, we transform it into a QUBO form. Employing a multi-level graph coarsening and refinement approach, we reduce the problem size to a range manageable by quantum computers, successfully utilizing CIMs to solve for the neural network’s global Lipschitz constant.
In this study, we focus on the FGL under $\ell_{\infty}$ perturbations, which approximates the maximum norm of the gradient operator and assumes that all activation patterns in the hidden layers are independent and possible. FGL serves as an upper bound for the exact global Lipschitz constant and has been utilized in previous works  (Raghunathan et al., 2018; Fazlyab et al., 2019; Latorre et al., 2020; Wang et al., 2022).
By exploiting the equivalence relation between FGL and the cut norm problem, our proposed HiQ-Lip method targets the estimation of the FGL which is the upper bound of $\ell_{\infty}$ global Lipschitz constant for two-layer fully connected neural networks. Initially, we convert the Lipschitz constant estimation of the neural network into a cut-norm problem. Subsequently, we iteratively coarsen the graph structure until obtaining the coarsest graph, solve the cut-norm problem on this graph using CIMs, and finally refine the solution by mapping it back to the original graph. This process results in an accurate estimation of the neural network’s global Lipschitz constant.
@@
Simulation experiments on fully connected neural networks demonstrate that HiQ-Lip achieves performance comparable to existing state-of-the-art (SOTA) methods in estimating the $\ell_{\infty}$ global Lipschitz constant for two-layer networks, while exhibiting faster computation speeds. When extended to multi-layer fully connected neural networks, applying our approach to every two consecutive layers yields reasonable estimates within a valid range. The extension of HiQ-Lip provides tighter estimates compared to the naive upper bound approach (Weight-Matrix-Norm-Product, MP method), particularly excelling in shallower networks. Comparing SOTA methods, especially in terms of running time, HiQ-Lip and HiQ-Lip extensions show up to two times speedup on two-layer neural networks and up to one hundred times speedup on multi-layer neural networks.
Our method is heuristic in nature and aims to obtain very approximate solutions for FGL estimation, and in the experiments, we obtain very close results to the SOTA method Geolip, which validates the effectiveness of HiQ-Lip.
Our main contributions are as follows:
@@
  • Firstly Innovations of Quantum Computing to Neural Network Robustness Estimation: We theoretically demonstrate the potential of small-scale quantum computing devices, such as CIMs, in the domain of neural network robustness estimation. We emphasize the NP-hard nature of the global Lipschitz constant estimation problem and introduce a quantum-classical hybrid solving method that synergizes quantum and classical systems to overcome the limited number of qubits in current quantum devices.
@@
  • Hierarchical algorithm Framework based on QUBO: We develop a QUBO-based solving algorithm framework that adopts a hierarchical approach. By transforming the $\ell_{\infty}$ global Lipschitz constant estimation problem of two-layer fully connected neural networks into a cut-norm problem, we achieve the first instance of utilizing CIMs to solve the global Lipschitz constant of neural networks efficiently on small-scale quantum devices.
  • Extension to Multi-Layer Networks with Quantum Device: For multi-layer neural networks, we demonstrate that HiQ-Lip provides tighter estimates compared to the naive upper bound (Weight-Matrix-Norm-Product) approach. Extending our method to neural networks with depths ranging from 3 to 5 layers, we obtain more precise estimates than the naive upper bound, particularly excelling in shallower three-layer networks.
  • Simulation Experiment Validation: We validate the effectiveness of HiQ-Lip on fully connected feedforward neural networks. For two-layer fully connected neural networks with one hidden layer, using CIMs with a limited number of qubits, we successfully measured the global Lipschitz constant for hidden layer sizes ranging from 8 to 256, achieving better results than existing SOTA methods.
@@ 2 Preliminaries
The global Lipschitz constant $L$ measures how fast the function $f$changes over all input pairs $(x,y)$ and is a key metric to evaluate the robustness of neural networks.
@@ 3 Transforming Lipschitz Constant Estimation into QUBO Formulation
where $A=W^{T}\text{diag}(u)$ and $u$ represents the activation pattern. In this task, $q$ is $\infty$ but $1$ is also introduced as a dual auxiliary.
@@
We begin by leveraging the duality between the $\ell_{1}$ and $\ell_{\infty}$ norms, which is well-known in optimization contexts  (Johnson, 2006; Martin, 1998). Specifically, the $\ell_{1}$ norm of a vector $v\in\mathbb{R}^{n}$ can be expressed using the maximum inner product between $v$ and a binary vector $z$ from the set $\{-1,1\}^{n}$:
@@
Hamiltonian is the dependence of many quantum devices to solve problems  (Cerezo et al., 2021; Glos et al., 2022). CIM and other quantum devices estimate the eigenvalue of Hamiltonian based on its characteristics. Both CIM and quantum annealers aim at finding the minimum of the QUBO problem. Define the Hamiltonian pointing in the direction of quantum evolution:
@@
[…]
However, directly solving this Hamiltonian requires $O(n+m)$ qubits, exceeding the capacity of current quantum devices (about 100 qubits) for practical networks (e.g., $n+m>784$ for MNIST networks)  (Proctor et al., 2022; Kim et al., 2023; Klimov et al., 2024; Pelofske et al., 2023).
@@ 4 HiQ-Lip for Lipschitz Constant Estimation
In this section, we propose a method called HiQ-Lip that aims to utilize small-scale quantum computers to efficiently estimate $\ell_{\infty}$-FGL. Due to the computational complexity of direct estimation, we employ a hierarchical solution strategy that treats the weights of the neural network as the edge weights of the graph and the neurons as the nodes. Based on the Equation 5 and 6, we construct a weighted undirected graph $G$ whose weight matrix is expressed as follows. $A=W^{T}\text{diag}(u)$, where $W\in\mathbb{R}^{n\times m}$ is the weight matrix between the input layer and the hidden layer, $u\in\mathbb{R}^{m}$ is the per-class weight vector between the hidden layer and the output layer.
@@ 4.1 Coarsening Phase
After each coarsening step, the number of vertices decreases approximately logarithmically. The time complexity of the $i$-th coarsening step is $\mathcal{O}(N_{i}^{2})$, where $N_{i}$ is the number of nodes at level $i$. Consequently, the overall time complexity of the Coarsening Phase is $\mathcal{O}(N^{2}\log N)$, where $N=n+m$. A detailed analysis is provided in Appendix A.1.
@@ 4.2 Refinement Phase
The Graph Refinement Phase iteratively improves the solution by mapping it back to finer graph levels. The primary steps include initializing the solution, computing gains for each node, and performing local optimizations using quantum solvers. Computing the gain for all nodes incurs a time complexity of $\mathcal{O}(N^{2})$. The local optimization step involves solving smaller QUBO problems of size $K$ on a quantum device, which contributes $\mathcal{O}(K^{\alpha})$ to the complexity, with $\alpha$ being a small constant. Overall, the Graph Refinement Phase operates with a time complexity of $\mathcal{O}(N^{2})$. See the Appendix  A.2 for a detailed analysis.
@@ 4.3 Algorithm Overview
[…]
@@
1:  Input: Weight matrix $A$, initial graph $G_{0}$
[…]
@@
[…]
@@
The HiQ-Lip algorithm achieves an overall time complexity of $\mathcal{O}(N^{2}\log N)$. See the Appendix A.4 for details. The Coarsening Phase dominates the complexity with $\mathcal{O}(N^{2}\log N)$ due to the iterative merging of nodes and updating of the adjacency matrix at each hierarchical level. Noticed that the running time of quantum devices is generally considered to have a large speedup over classical computers (Mohseni et al., 2022; Avkhadiev et al., 2020; Di Meglio et al., 2024). Even for problems of exponential complexity, the running time of a quantum device can be considered a fraction of the task time on a classical computer. In the Refinement Phase, the complexity is $\mathcal{O}(N^{2})$, primarily from computing gains for node optimizations and solving smaller QUBO problems using quantum devices. Consequently, the combined phases ensure that HiQ-Lip scales efficiently for neural networks of moderate size, leveraging quantum acceleration to enhance performance without exceeding polynomial time bounds.
5 Extension to Multi-Layer Networks
Directly estimating the global Lipschitz constant for multi-layer networks is challenging due to the complexity of tensor cut-norm problems. We extend our method by approximating the network’s Lipschitz constant as the product of the Lipschitz constants of individual layer pairs. By leveraging the triangle inequality $\|A\|\|B\|\geq\|AB\|$, we transform the upper bound estimation of the FGL into a layer-wise product of weight norms. This approach simplifies the estimation process by decomposing the multi-layer problem into manageable two-layer subnetworks, thereby enabling more efficient and scalable computations. It is inspired by the work of  (Leino et al., 2021b; Bartlett et al., 2017; Szegedy et al., 2014).
Our approach involves decomposing the multi-layer network into a series of two-layer subnetworks and estimating the Lipschitz constant for each subnetwork. We then combine these estimates to obtain the overall Lipschitz constant of the network. The key steps are as follows:
  1. 1.
    Processing the Output Layer: We start by focusing on a specific output neuron (e.g., corresponding to a particular class in classification tasks). We extract the weights associated with this output neuron and incorporate them into the previous layer’s weights.
  2. 2.
    Constructing Adjacency Matrices: For layers $l$ and $l+1$, build the corresponding graph model. We construct an adjacency matrix based on the modified weight matrices. These adjacency matrices represent the connections between neurons in the two layers.
  3. 3.
    Computing Cut-Norms via HiQ-Lip: We formulate a QUBO problem for each adjacency matrix and use the hierarchical method to estimate $\|A^{l}\|_{\infty\to 1}$. This yields an estimate of the cut-norm for each layer pair.
  4. 4.
    Combining Results: We multiply the cut-norm estimates for all layer pairs and adjust for scaling factors to compute the overall Lipschitz constant estimate.
The coefficient $\frac{1}{2^{d-2}}$ of FGL in a multi-layer neural network is used by  (Latorre et al., 2020). For the $d$-layer network, we decompose the estimate as:
-$\ell_{\infty}{\text{-FGL}}\leq\max\left\{\frac{1}{2^{d-2}}\prod_{l=1}^{d-1}\|A^{l}\|_{\infty\to 1}\right\},$ (17)
where $A^{l}$ is the processed weight matrix between layers $l$ and $l+1$, and $A=W^{T}\text{diag}(u)$. We refer to methods with such coefficients as HiQ-Lip MP A.
However, we found in our experiments that this approach gives loose estimates compared with GeoLip. To obtain more robust estimates for 4 and 5 layers networks, we propose a new coefficient as $\frac{1}{2^{d-2}}\frac{1}{d^{d-3}}$ to estimate FGL for fully connected neural networks with 3 to 5 layers. We refer to methods with such coefficients as HiQ-Lip MP B. The approach of using a more compact coefficient has been applied in previous work  (Bartlett et al., 2017).
Intuitively, the new coefficients are introduced to make the estimation of the Lipschitz constant more compact in multi-layer neural networks. According to the triangle inequality, if the independence assumption between the layers does not hold perfectly, the original product estimate will appear loose. Therefore, additional correction coefficients are introduced to weaken the influence of each layer contribution. It can be regarded as a regularization based on the number of network layers to the original product upper bound. Especially in deep networks, the nonlinear complexity makes the product grow too fast, resulting in the estimated upper bound is much larger than the actual value.
These approaches leverage the sub-multiplicative property of norms and provide a tighter upper bound compared to naive methods.

### v2

# HiQ-Lip: A Hierarchical Quantum-Classical Method for Global Lipschitz Constant Estimation of ReLU Networks

@@ Abstract
Estimating the global Lipschitz constant of neural networks is crucial for understanding and improving their robustness and generalization capabilities. However, precise calculations are NP-hard, and current semidefinite programming (SDP) methods face challenges such as high memory usage and slow processing speeds. In this paper, we propose HiQ-Lip, a hybrid quantum-classical hierarchical method that leverages quantum computing to estimate the global Lipschitz constant. We tackle the estimation by converting it into a Quadratic Unconstrained Binary Optimization problem and implement a multilevel graph coarsening and refinement strategy to adapt to the constraints of contemporary quantum hardware. Our experimental evaluations on fully connected neural networks demonstrate that HiQ-Lip not only provides estimates comparable to state-of-the-art methods but also significantly accelerates the computation process. In specific tests involving two-layer neural networks with 256 hidden neurons, HiQ-Lip doubles the solving speed and offers more accurate upper bounds than the existing best method, LiPopt. These findings highlight the promising utility of small-scale quantum devices in advancing the estimation of neural network robustness.
@@ 1 Introduction
Neural networks have achieved remarkable success in various fields such as computer vision, natural language processing, and autonomous driving, establishing themselves as core technologies in modern artificial intelligence (Zhao et al. 2024; Forner and Ozcan 2023; Paniego et al. 2023). Despite these advancements, the robustness and generalization capabilities of neural networks remain active areas of research (Djolonga et al. 2021; Bennouna et al. 2023). Research on adversarial security in neural networks often starts from the perspective of non-robustness (Weng et al. 2025). The global Lipschitz constant is a critical metric for measuring the robustness of a neural network’s output to input perturbations, playing a significant role in understanding and enhancing model robustness (Leino et al. 2021).
@@
However, accurately computing the global Lipschitz constant is an NP-hard problem, particularly for deep networks with nonlinear activation functions like ReLU (Jordan and Dimakis 2020). Due to this computational challenge, researchers have developed various approximation methods to estimate upper bounds of the global Lipschitz constant. Among these, the Formal Global Lipschitz constant (FGL) assumes that all activation patterns in the hidden layers are independent and possible, thereby providing an upper bound for the exact global Lipschitz constant (Szegedy et al. 2013; Virmaux and Scaman 2018).
@@
In recent years, advancements in quantum computing have offered new avenues for tackling NP-hard problems (Khumalo et al. 2022; Choi 2008; Chatterjee et al. 2024). Given the computational complexity of global Lipschitz constant estimation, quantum computation is seen as a potential solution. Specifically, quantum devices based on the Quadratic Unconstrained Binary Optimization (QUBO) model, such as Coherent Ising Machines (CIMs) or quantum annealers, have demonstrated unique potential in solving complex combinatorial optimization problems (Inagaki et al. 2016; Date et al. 2019). However, the limited number of qubits in current quantum computers poses significant challenges when directly applying them to neural network robustness evaluations. Although numerous works have attempted to address large-scale QUBO problems by employing strategies like divide-and-conquer and QUBO formula simplification (Pelofske et al. 2021; Zhou et al. 2023), applications in neural network robustness assessment remain largely unexplored.
@@
To bridge this gap and harness the acceleration capabilities of quantum computing for FGL estimation of neural networks, we propose a quantum-classical hybrid hierarchical solving method named HiQ-Lip. HiQ means a multilevel solution method for quantum-classical mixing, while Lip stands for global Lipschitz constant. By reformulating the global Lipschitz constant estimation problem as a cut-norm problem, we transform it into a QUBO form. Employing a multilevel graph coarsening and refinement approach, we reduce the problem size to a range manageable by quantum computers, successfully utilizing CIMs to solve for the neural network’s global Lipschitz constant.
@@
Our method is heuristic in nature and aims to obtain approximate solutions for FGL estimation. Simulation experiments on fully connected neural networks demonstrate that HiQ-Lip achieves performance comparable to existing state-of-the-art methods in estimating the $\ell_{\infty}$ global Lipschitz constant for two-layer networks, while exhibiting significantly faster computation speeds. When extended to multi-layer fully connected neural networks, our approach yields tighter estimates compared to the naive upper bound approach (Weight-Matrix-Norm-Product), particularly excelling in shallower networks. Comparing SOTA methods, especially in terms of running time, HiQ-Lip shows substantial acceleration, achieving up to two times speedup on two-layer neural networks and up to one hundred times speedup on multi-layer neural networks.
The main contributions of this paper are summarized as follows:
@@
  • First Innovations of Quantum Computing to Neural Network Robustness Estimation: We theoretically demonstrate the potential of small-scale quantum computing devices, such as CIMs, in neural network robustness estimation and propose a quantum-classical hybrid method to address limitations posed by the current quantum devices’ qubit capacity.
@@
  • Hierarchical Algorithm Framework Based on QUBO: We develop a QUBO-based hierarchical solving algorithm framework, reformulating the $\ell_{\infty}$ global Lipschitz constant estimation for two-layer fully connected neural networks as a cut-norm problem, and employing CIMs to solve this efficiently.
  • Extension to Multi-Layer Networks: For multi-layer networks, HiQ-Lip provides tighter estimates compared to existing naive approaches, demonstrating its effectiveness particularly for networks with depths ranging from three to five layers.
@@ 2 Preliminaries
The global Lipschitz constant $L$ measures how fast the function ffchanges over all input pairs $(x,y)$ and is a key metric to evaluate the robustness of neural networks.
@@ 3 Transforming Lipschitz Constant Estimation into QUBO Formulation
where $A=W^{T}\text{diag}(u)$ and $u$ represent the activation pattern. In this task, $q$ is $\infty$ but $1$ is also introduced as a dual auxiliary.
@@
We begin by leveraging the duality between the $\ell_{1}$ and $\ell_{\infty}$ norms, which is well-known in optimization contexts  (Johnson 2006; Martin 1998). Specifically, the $\ell_{1}$ norm of a vector $v\in\mathbb{R}^{n}$ can be expressed using the maximum inner product between $v$ and a binary vector $z$ from the set $\{-1,1\}^{n}$:
@@
Many quantum optimizers minimize an Ising-form Hamiltonian that encodes the objective  (Cerezo et al. 2021; Glos et al. 2022). CIM and other quantum devices seek low-energy configurations of Hamiltonian based on its characteristics. Both CIM and quantum annealers aim at finding the minimum of the QUBO problem. Define the Hamiltonian pointing in the direction of quantum evolution:
@@
the problem of estimating -$\ell_{\infty}{\text{-FGL}}$ becomes minimizing $H$, which is a QUBO problem. Then the CIM or other quantum device could solve it by the QUBO formulation.
However, directly solving this Hamiltonian requires $O(n+m)$ qubits, exceeding the capacity of the quantum devices available (about 100 qubits) when this study was conducted for practical networks (e.g., $n+m>784$ for MNIST networks)  (Proctor et al. 2022; Kim et al. 2023; Klimov et al. 2024; Pelofske et al. 2023).
@@ 4 HiQ-Lip for Lipschitz Constant Estimation
In this section, we propose a method called HiQ-Lip that aims to utilize small-scale quantum computers to efficiently estimate $\ell_{\infty}$-FGL. Due to the computational complexity of direct estimation, we employ a hierarchical solution strategy that treats the weights of the neural network as the edge weights of the graph and the neurons as the nodes. Based on the equations 5 and 6, we construct a weighted undirected graph $G$ whose weight matrix is expressed as follows. $A=W^{T}\text{diag}(u)$, where $W\in\mathbb{R}^{n\times m}$ is the weight matrix between the input layer and the hidden layer, $u\in\mathbb{R}^{m}$ is the per-class weight vector between the hidden layer and the output layer.
@@ 4.1 Coarsening Phase
After each coarsening step, the number of vertices decreases approximately logarithmically. The time complexity of the $i$-th coarsening step is $\mathcal{O}(N_{i}^{2})$, where $N_{i}$ is the number of nodes at level $i$. Consequently, the overall time complexity of the Coarsening Phase is $\mathcal{O}(N^{2}\log N)$, where $N=n+m$.
@@ 4.2 Refinement Phase
The Graph Refinement Phase iteratively improves the solution by mapping it back to finer graph levels. The primary steps include initializing the solution, computing gains for each node, and performing local optimizations using quantum solvers. Computing the gain for all nodes incurs a time complexity of $\mathcal{O}(N^{2})$. The local optimization step involves solving smaller QUBO problems of size $K$ on a quantum device, which contributes $\mathcal{O}(K^{\alpha})$ to the complexity, with $\alpha$ being a small constant. Overall, the Graph Refinement Phase operates with a time complexity of $\mathcal{O}(N^{2})$.
@@ 4.3 Algorithm Overview
As shown in Algorithm 1, HiQ-Lip begins with the initial graph $G_{0}$ and, through lines 3 to 9, executes the coarsening phase by iteratively merging nodes to produce progressively smaller graphs suitable for quantum processing. In line 10, it leverages quantum acceleration by solving the resulting QUBO problem on the coarsest graph using CIM or other quantum devices to efficiently obtain an initial solution. Finally, lines 11 to 15 implement the refinement phase, where the algorithm maps this solution back onto finer graph layers and performs local optimizations to accurately estimate the -$\ell_{\infty}\text{-FGL}$.
@@
1:Input: Weight matrix AA, initial graph $G_{0}$
2:Output: Estimated Lipschitz constant -$\ell_{\infty}\text{-FGL}$
@@
16: Compute -$\ell_{\infty}\text{-FGL}$ from the final solution
@@
The HiQ-Lip algorithm achieves an overall time complexity of $\mathcal{O}(N^{2}\log N)$. The Coarsening Phase dominates the complexity with $\mathcal{O}(N^{2}\log N)$ due to the iterative merging of nodes and updating of the adjacency matrix at each hierarchical level. Noticed that the running time of quantum devices is generally considered to have a large speedup over classical computers (Mohseni et al. 2022; Avkhadiev et al. 2020; Di Meglio et al. 2024). Even for problems of exponential complexity, the running time of a quantum device can be considered a fraction of the task time on a classical computer. In the Refinement Phase, the complexity is $\mathcal{O}(N^{2})$, primarily from computing gains for node optimizations and solving smaller QUBO problems using quantum devices. Consequently, the combined phases ensure that HiQ-Lip scales efficiently for neural networks of moderate size, leveraging quantum acceleration to enhance performance without exceeding polynomial time bounds.
5 Extension to Deep $L$-Layer ReLU Networks
Estimating the global $\ell_{\infty}\!\to\!\ell_{1}$ Lipschitz constant of deep ReLU networks is NP‑hard because the Jacobian $J(x)=W_{L}D_{L-1}\!\cdots D_{1}W_{1}$ is a (L1)(L{-}1)-multilinear function of the binary activation patterns $D_{\ell}=\mathrm{diag}(s_{\ell}),\;s_{\ell}\!\in\!\{0,1\}^{h_{\ell}}$. A naive polynomial–to‑quadratic reduction would create $\Theta\!\bigl((\sum_{\ell}h_{\ell})^{2}\bigr)$ ancilla variables, rendering quantum‑classical QUBO solvers impractical beyond two or three layers.
5.1 Layer‑wise Max‑Cut recursion
Inspired by the cut‑norm formulation of two‑layer networks (Wang et al. 2022; Latorre et al. 2020), we retain a quadratic objective by unfolding the product one layer at a time:
$\|P_{k}\|_{\infty\!\to 1}$ $=\max_{x\in\{\pm 1\}^{n_{k}},\;y\in\{\pm 1\}^{m_{k}}}x^{\!\top}P_{k}\,y,$
$P_{k}$ $=W_{k+1}D_{k}P_{k-1}$
with $P_{0}=W_{1}$. Each subproblem is a weighted MaxCut that admits the Goemans–Williamson SDP approximation factor $0.878$ (Goemans and Williamson 1995). The global constant is then
$L_{\infty\!\to 1}~=~\max_{s_{1},\dots,s_{L-1}}\bigl\|P_{L-1}(s_{1{:}L-1})\bigr\|_{\infty\!\to 1}.$ (17)
Error accumulation.
If the cut‑norm optimiser at depth $k$ yields $(1+\varepsilon_{k})$ approximation error, we obtain the telescoping bound $L_{\infty\!\to 1}\;\leq\;\bigl(\prod_{k=1}^{L-1}(1+\varepsilon_{k})\bigr)\widehat{L}_{\infty\!\to 1},$ where $\widehat{L}_{\infty\!\to 1}$ is the estimate returned by the recursion. In practice $\varepsilon_{k}$ stays under $10^{-2}$ for $h_{k}\!\leq\!256$, yielding a depth‑linear error factor that is comparable to the chordal‑sparsity SDP of Xue et al. (2022); Wang et al. (2024).
5.2 Block‑wise product upper bound
For very deep or convolutional architectures, we adopt the block‑product strategy: split the network into $B$ contiguous blocks of at most $b$ layers ($b=2$ or $3$ works well empirically), estimate a tight constant $\gamma_{i}$ for each block via the above SDP/QUBO routine, then combine them multiplicatively:
$L_{\infty\!\to 1}\;\;\leq\;\;\frac{1}{c_{b}^{\,L-B}}\prod_{i=1}^{B}\gamma_{i},\qquad c_{b}=2^{\,b-1}.$ (18)
Inequality (18) follows from the sub‑multiplicativity of $\|\cdot\|_{\infty\!\to 1}$ and the triangle inequality $\|A\|\|B\|\!\geq\!\|AB\|$.
5.3 Depth‑adaptive damping coefficients
Latorre et al. (2020) use the global factor $2^{-(L-2)}$ for fully connected nets, which becomes loose for L>4L\!>\!4. Motivated by Bartlett–Foster–Telgarsky’s exponential depth dependence of margin bounds (Bartlett et al. 2017, Thm. 3.3), we introduce a depth‑adaptive coefficient $c_{L}=2^{\,L-2}L^{\,L-3},$ leading to the estimator
$\widehat{L}_{\infty\!\to 1}^{(\mathrm{HiQ\text{-}Lip\;MP\text{-}B})}~=~\frac{1}{c_{L}}\;\prod_{\ell=1}^{L-1}\!\|A^{\ell}\|_{\infty\!\to 1},\quad A^{\ell}=W_{\ell+1}D_{\ell}.$ (19)
On CIFAR‑10 MLPs with L=8L\!=\!8, this coefficient tightens the GeoLip upper bound by $15\%$ while remaining GPU‑friendly.
5.4 Discussion
Relation to prior work.
Our framework generalises GeoLip(Wang et al. 2022) ($b=L$), integrates the spectral‑product heuristics of Fazlyab et al. (2019), and complements chordal‑SDP scaling (Xue et al. 2022; Wang et al. 2024) by targeting the non‑spectral cut‑norm. The depth‑adaptive coefficient is conceptually similar to the layer‑wise contraction in Araujo et al. (2023), but derived for $\ell_{\infty}$ perturbations.